# Nuclear Cold QCD: Review and Future Strategy

### v1

@@ VII.2 Ultraperipheral Heavy-Ion Collisions at the LHC (2026–2035)
In the coming years, new measurements of open heavy-flavor hadrons, quarkonia, and jets in ultraperipheral heavy-ion collisions will provide high accuracy constraints on quark and gluon PDFs at low $x$ that are unaffected by final-state interactions. This program will benefit from the high center-of-mass energies at the LHC. The LHC’s high-luminosity upgrade for Run 4 (HL-LHC) starting in 2030 will enhance precision and detector coverage, leading to stronger constraints on PDFs in a wider range of $x$ and $Q^{2}$. By combining the low $x$ reach of UPC measurements with the precision of $e+A$ collisions at the EIC, we the kinematic reach of the combined results will place tighter constraints on nPDFs.
@@ VII.3 Fixed-Target Program at the EIC
The Long Range Plan for Nuclear Sience prioritizes the EIC as the highest priority new US facility [493]. The EIC will collide electron beams of up to 18 GeV with ion beams of up to 100 GeV/nucleon [494]. The RHIC facility will be shut down and the EIC will be constructed in its stead. This transition will eliminate studies of $p$+$p$, $p+A$, and $A+A$ collisions in the US at RHIC, making the LHC the only facility worldwide for ultrarelativistic hadronic collisions.
@@
Figure 20: Conceptual diagram of the EIC collision region with fixed target. Collider $e+A$ collisions will produce particles across all angles, while in the $A$-going direction (left) particles from fixed target $A+A^{\prime}$ collisions can be measured, and in the $e$-going direction (right) lower energy $e+A^{\prime}$ collisions can be studied. This configuration woud be equivalent to running three experiments simultaneously.
@@ VII.4 The Forward Physics Facility
  • searches for new physics beyond the Standard Model, including light weakly-interacting particles including dark photons, axion-like particles, and sterile neutrinos;
@@
Most relevant to these studies, the FPF neutrino studies as a fixed-target experiment could be complementary to EIC program. The energy flux of neutrinos detected in the FPF, peaked around 1 TeV, is very broad. In a fixed target configuration, this energy flux would correspond to $\sqrt{s_{NN}}\approx 45^{+45}_{-23}$ GeV due to energy spread of the neutrino flux. Such a fixed-target configuration would nicely complement the EIC physics program with $\nu+A$ DIS measurements at energies similar to those of the EIC. For example, studies of coherent power corrections predicts a measurable difference in the shadowing pattern of the $F_{2}^{A}$ and $F_{3}^{A}$ [150] structure functions as well as significant low- and moderate-$Q^{2}$ modifications of QCD sum rules based on these structure functions [506]. These complementary measurements could help determine the origin of shadowing.
@@ VII.5 The ideal experiment
Surveying the experimental landscape for future studies of cold nuclear matter (CNM) effects, one can envision an ideal setup that combines both collider and fixed-target modes to maximize physics reach. In collider mode, a proton beam could interact with various nuclear beams at $\sqrt{s_{NN}}\sim 100$ GeV. This configuration provides access to re;atove;u high $p_{T}$ and broad $x_{F}$ coverage at both forward and backward of the center of the collision, enabling sensitive probes of CNM dynamics.
In fixed-target mode, the same proton beam could interact with an internal gas target, as pioneered by the SMOG2 system at LHCb [489, 490]. Realistic gas targets include H, D, He, Ne, Ar, Kr, and Xe. Solid nuclear targets such as Au or Pb could also be considered. Such configurations would result in a center-of-mass energy of $\sqrt{s_{NN}}\sim 10$ GeV. While lower in energy, the fixed-target mode offers high luminosity and access to large $x_{F}$, which are essential for highlighting the strong suppression pattern in that region.
High-precision silicon vertex detectors would be essential to (i) improve vertex resolution, which is particularly challenging with solid targets, and (ii) reconstruct heavy-flavor hadrons and quarkonia, providing complementary observables to probe both initial- and final-state effects.

### v2

@@ VII.2 Ultraperipheral Heavy-Ion Collisions at the LHC (2026–2035)
In the coming years, new measurements of open heavy-flavor hadrons, quarkonia, and jets in ultraperipheral heavy-ion collisions will provide high accuracy constraints on quark and gluon PDFs at low $x$ that are unaffected by final-state interactions. This program will benefit from the high center-of-mass energies at the LHC. The LHC’s high-luminosity upgrade for Run 4 (HL-LHC) starting in 2030 will enhance precision and expand detector coverage, leading to stronger constraints on PDFs in a wider range of $x$ and $Q^{2}$. By combining the low $x$ reach of UPC measurements with the precision of $e+A$ collisions at the EIC, the kinematic reach of the combined results will place tighter constraints on nPDFs.
@@ VII.3 Fixed-Target Program at the EIC
The Long Range Plan for Nuclear Sience listed the EIC as the highest priority new US facility [497]. The EIC will collide electron beams of up to 18 GeV with ion beams of up to 100 GeV/nucleon [498]. The RHIC facility will be shut down and the EIC will be constructed in its stead. This transition will eliminate studies of $p$+$p$, $p+A$, and $A+A$ collisions in the US at RHIC, making the LHC the only facility worldwide for ultrarelativistic hadronic collisions.
@@
Figure 20: Conceptual diagram of the EIC collision region with a fixed target. Collider $e+A$ collisions will produce particles across all angles, while in the $A$-going direction (left) particles from fixed target $A+A^{\prime}$ collisions can be measured, and in the $e$-going direction (right) lower energy $e+A^{\prime}$ collisions can be studied. This configuration woud be equivalent to running three experiments simultaneously.
@@ VII.4 The Forward Physics Facility
  • searches for new physics beyond the Standard Model, including light weakly-interacting particles such as dark photons, axion-like particles, and sterile neutrinos;
@@
Most relevant to these studies, the FPF neutrino studies as a fixed-target experiment could be complementary to EIC program. The energy flux of neutrinos detected in the FPF, with a peak around 1 TeV, is very broad. In a fixed target configuration, this energy flux would correspond to $\sqrt{s_{NN}}\approx 45^{+45}_{-23}$ GeV due to the energy spread of the neutrino flux. Such a fixed-target configuration would nicely complement the EIC physics program with $\nu+A$ DIS measurements at energies similar to those of the EIC. For example, studies of coherent power corrections predicts a measurable difference in the shadowing pattern of the $F_{2}^{A}$ and $F_{3}^{A}$ [150] structure functions as well as significant low- and moderate-$Q^{2}$ modifications of QCD sum rules based on these structure functions [510]. These complementary measurements could help determine the origin of shadowing.
@@ VII.5 The ideal experiment
Surveying the experimental landscape for future studies of CNM effects, one can envision an ideal setup that combines both collider and fixed-target modes to maximize physics reach. In collider mode, a proton beam could interact with various nuclear beams at $\sqrt{s_{NN}}\sim 100$ GeV. This configuration provides access to relatively high $p_{T}$ and broad $x_{F}$ coverage both forward and backward of the center of the collision, enabling sensitive probes of CNM dynamics.
In fixed-target mode, the same proton beam could interact with an internal gas target, as pioneered by the SMOG2 system at LHCb [493, 494]. Realistic gas targets include H, D, He, Ne, Ar, Kr, and Xe. Solid nuclear targets such as Au or Pb could also be considered. Such configurations would result in a center-of-mass energy of $\sqrt{s_{NN}}\sim 10$ GeV. While lower in energy, the fixed-target mode offers high luminosity and access to large $x_{F}$, both of which are essential for highlighting the strong suppression pattern in that region.
High-precision silicon vertex detectors would be essential to improve vertex resolution, particularly challenging with solid targets, and reconstruct heavy-flavor hadrons and quarkonia, providing complementary observables to probe both initial- and final-state effects.